OpenStudy (anonymous):

Find the difference 7 & 7/12-3 & 5/9

OpenStudy (anonymous):

I keep getting things like 0.3518518518518519

OpenStudy (anonymous):

That's what happens when you use a calculator :)

OpenStudy (anonymous):

yeah! Lol.

OpenStudy (anonymous):

The LCD is 36, right?

OpenStudy (anonymous):

Yup....i think so lol

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

\[7\frac{7}{12}=7\frac{21}{36}\] \[3\frac{5}{9}=3\frac{20}{36}\]

OpenStudy (anonymous):

You see how we got to this?

OpenStudy (anonymous):

umm.....

OpenStudy (anonymous):

jw is the answer 4 12/21 ...or somthing..i think.. O.O

OpenStudy (anonymous):

\[4\frac{1}{36}\]

OpenStudy (anonymous):

:O how did you get that....i tried to cross multiply like it was said on the last thingy!!!

OpenStudy (anonymous):

No, they were wrong hehe. You don't cross multiply

OpenStudy (anonymous):

You find the LCD and rewrite the fraction in terms of the LCD

OpenStudy (anonymous):

:O ohh no lol. So your right! :P